OpenStudy (anonymous):

a culture started with 6000 bacteria. after 4 hours it grew to 6600. predict how many bacteria will be present after 10 hours

OpenStudy (anonymous):

From 6000 to 6600 it grew by 600, right?

OpenStudy (anonymous):

calculate per hour growth

OpenStudy (anonymous):

So to find how much it grows in one hour you would divide 600 / 4 = 150

OpenStudy (anonymous):

Then multiply by 10, so 150 * 10 = 1500.

OpenStudy (anonymous):

growth per hour in 4 hour=(6600-6000)/4=150

OpenStudy (anonymous):

@uselly To give medals click "Best Response".